OpenStudy (mackenzie2013):

If x is to 9 as 8 is to 12, then x is equal to ?

OpenStudy (mertsj):

cross multiply

OpenStudy (anonymous):

\[\frac{ x }{ 9 }=\frac{ 8 }{ 12 }\]

OpenStudy (anonymous):

12x=72 12x/12=x 72/12=6 x=6

OpenStudy (mackenzie2013):

Thanks ! :)

OpenStudy (anonymous):

You're Welcome